OpenStudy (sphott51):

Simplify the expression (exponential)... (m^-6 n^-3)/ (m^-13 n^-1)

Nnesha (nnesha):

when we multiply same bases we should `add` the exponents \[\huge\rm x^m \times x^n=x^{m+n}\] and when divide same base , `subtract` the exponents \[\huge\rm \frac{ x^m }{ x^n }=x^{m-n}\]

Nnesha (nnesha):

\[\large\rm x^{-m}= \frac{1}{x^m}\] flip the fraction to convert negative to positive exponent

OpenStudy (sphott51):